i had to solve the equation for y. the equation was (sqrt of 2y) +3=11, my answer was 4 ? please help.i hope i typed the equation right the 2y are under the radical sign.
Found 2 solutions by stanbon, ikdeep:
Answer by stanbon(57387)

sqrt(2y) + 3 = 11
sqrt(2y)=8
Square both sides to get:
2y=64
y=32
Check the answer:
sqrt(2(32))+3=11
sqrt(64)+3+11
8+3=11
11=11
Cheers,
Stan H.

Answer by ikdeep(226)

(sqrt of 2y) +3=11
first subtract both sides by 3 and we get ..
(sqrt of 2y) = 8
taking square of both sides we get
2y = 64
divide both sides by 2 and we get..
y = 32

Now if you want to verify you answer ,,,you can put the values of y in the given equation and if you get LHS = RHS,,,this means that you answer is correct..
hope this will help you
Please feel free to revert back for any further queries.